# Quantum-enhanced Markov Chain Monte Carlo for systems larger than your Quantum Computer

### v1

@@ 4.1 Naive CG - local groups
Ignoring the surrounding spins is an approximation that understandably has limitations and specifically raises two main issues. Firstly, the Hamming distance of proposal steps is fundamentally restricted to the size of the group. While this is better than classical local move approaches, it still puts a bound on the jumps that the proposal steps can make, and is more likely to get stuck in local minima (as with the classical local proposal) and partially cancels one of the key properties that lead to the success of the QeMCMC - namely the large Hamming distance update proposals.
The second issue is that the new problem Hamiltonian $H_{prob}^{\prime}$ is missing contributions from the ignored spins making the evolution fundamentally incomplete. This partly cancels the other advantage of QeMCMC highlighted earlier, that it includes the problem’s Hamiltonian. If one were to consider a 1D Ising model where only neighboring spins interact, this choice of CG is likely to perform adequately, as the boundary interaction terms of the group is constant (2) with cluster size, meaning the impact of missing contributions from ignored spins scales favourably with group size. The same is not true for general Ising models with higher connectivity.
@@
4.2 Improved local groups
@@
Explicitly, the interaction matrix is the same as in the local grouping, $\tilde{J}=J_{i\in g_{l},j\in g_{l}}$, but the non-selected spins are treated as environment, adding them to the field of the $i^{th}$ selected spin as:
$\tilde{h}_{i}=h_{i}+\sum_{j\notin g_{l}}^{n}J_{ij}s_{j}$ (14)
In this way, the accuracy of the Hamiltonian is increased from the local group method without the addition of any quantum gates. The effective Hamiltonian $H_{prob}^{\prime}$ now fully represents the relative energy of any two states of the group. As will become clear in Section 5, improved local grouping refines the naive local groups, however there is still tail-off for small $q/n$.
@@ 4.3 Multiple groups
// Initialisation
2 $\boldsymbol{s}\leftarrow$ initial spin configuration;
@@ 8      $index\leftarrow\text{spins in }i^{th}\text{ group}$;
9      $s_{g}\leftarrow\text{subset of }\boldsymbol{s}\text{ using }index$;
@@ 10      $s_{g}^{\prime}\leftarrow$ QeMCMC ($s_{g}$);
      // update $\boldsymbol{s}^{\prime}$ with results from $i^{th}$ group’s QeMCMC
11      $\boldsymbol{s}^{\prime}[index]\leftarrow s_{g}^{\prime}$ ;
@@
5 Numerical Results
In this section we analyse the results of our numerical experiments of Ising model CGeMCMC, while further details of the implementation can be found in A.
5.1 Spectral gap
To quantify the CGQeMCMC, the effect of reducing $q$ on the convergence of the resulting Markov chain must be understood, with the particular aim of understanding the different coarse grained proposals at $q=\sqrt{n}$. This is done by evaluating the spectral gap of the CG methods for randomly initialised fully-connected Ising models in the range $4-10$ spins. Figure 1 shows the decrease of spectral gap with $q/n$ for each of the CG methods and $n=\{6,8,10\}$. Clearly the decay is worse for techniques that do not multiple sample, suggesting that the locality of low $q$ restricts the proposal. This effect is dependent on system size, however it is difficult to quantify using only the small systems at our disposal. There is also a temperature dependence, as the performance of small $q$ is inextricably linked with the performance of the classical local proposal. For analysis of temperature dependence for a 9 spin instance see Figure 2.
@@ 5.1 Spectral gap
In analogy with [1] Figure 2a, the different update proposal strategies are compared for a representative set of $100$ randomly initialised $9$ spin fully-connected Ising model instances in Figure 2. The QeMCMC (blue) is far superior to any classical strategy (orange and green) in the low temperature regime. The CGQeMCMC with $3$ groups of $3$ qubits (red) is also superior to the classical strategy, however the speedup is clearly dampened. If one evaluates a single $3$ spin group on the lattice each quantum proposal (transparent blue) the quantum speedup is lost for most temperature ranges for this system size.
Figure 2: Spectral gap varying with temperature for $100$ randomly initialised $9$ spin fully-connected Ising model. The classical uniform (orange) and local (green) strategies are clearly improved at low temperatures by the QeMCMC strategy (blue) [1]. The red data represents CGQeMCMC with $q=\sqrt{n}=3=n_{g}$ ($3$ groups of $3$ spins). The light blue data is $q=\sqrt{n}$ but $n_{g}=1$ ($1$ group of $3$ spins).
There is evident promise of quantum speedup using merely $\sqrt{n}$ qubits, however speedup at a fixed system size is less important. The scaling behaviour of the proposal step is necessarily analysed at $T=1$ in Figure 3. Equation 15 is the exponential fitting function used in the figures, where a small $k$ represents good scaling of a proposal strategy. The data points for larger systems are less reliable due to a much restricted sample size $(50)$ for the average spectral gap. This leads to added uncertainty in the fits, and increased reliance on the convergence analysis in Section 5.2.
@@
Figure 3: Spectral gap varying with problem size for $500$ randomly initialised fully-connected Ising models of each integer number of spins. Only $50$ systems of $9$ and $10$ spins for CG techniques were evaluated due to computational limitations. Each data set has least squares exponential fits with fitting function Equation 15. As the system size increases, the spectral gap of classical uniform (orange) and local (green) strategies clearly decrease faster than those of the QeMCMC (blue) and CGQeMCMC (red, light blue) proposals. For a given proposal, the value of $k$ determined by best fit is as follows: Uniform: $0.97(1)$, Local $0.92(4)$, QeMCMC: $0.335(8)$, local group: $0.77(2)$, Improved local group: $0.64(2)$, multiple sample: $0.50(3)$. $k$ values for $T=0.1$ and $T=10$ are listed in appendix 1 alongside the appropriate figures.
@@
For the multiple sample proposal strategy, we see a quantum enhancement factor of $k_{QEF}=1.84$ against the best classical approach - the local proposal. In this data, the QeMCMC gains an enhancement of $2.7$, a slightly reduced value than reported in [1] for the same temperature. Figure 3 depends largely on the temperature while Figure 2 depends on the problem size. B provides multiple figures that explores these relations. The corresponding quantum speedups in $k$ for $T=0.1$ and $T=10$ are $2.17$ and $2.6$ respectively. Of course, the problem instance type (1D, 2D, fully connected, sparse etc.) influences the outcome of these experiments, however these topics are beyond the scope of this work.
@@ 5.2 Convergence analysis
Markov chains using local, uniform and CGQeMCMC are used to evaluate the average energy and magnetisation of the stationary distribution of exemplar fully-connected Ising models. Systems of size $16$, $25$ and $36$ (see B for $15$ and $36$ spin results) are the focus so that $\sqrt{n}$ qubits can be easily compared with classical methods. The stationary distribution of the $25$ spin instance at $T=1$ is primarily composed of just a few low energy states. The Boltzmann probabilities of the ground state and $1^{st}$ $3^{rd}$ excited states are: $68.6\%$, $25.4\%$, $4.2\%$, $0.5\%$. If one were to sample lower temperature distributions the problem essentially becomes one of optimisation, while at higher energy the problem becomes less difficult as one transitions from the ferromagnetic phase. See B for temperature and system size analysis.
@@
Figure 4: Average energy of $10$ chains for uniform (orange), local(green), Improved local group (light blue) and multiple sample (red) with $q=\sqrt{n}=5=n_{g}$ ($5$ groups of $5$ spins) for $T=1$. Right figure is close-up of the low energy region. The classical local proposal has a large offset from the true average energy (black line). This is greatly reduced by CGQeMCMC. The uniform proposal suffers from extremely slow convergence but does not get stuck in local minima.
@@
Figure 4 shows the effective convergence of CGQeMCMC, however it is clear that the problem of getting stuck in local minima is inherited as the average energy over $10$ Markov chains is biased by a few that have not found the minimum state. One can crudely evaluate convergence in this relatively low temperature regime by calculating how many chains have discovered the minimum state. In this example, of the $10$ chains for each proposal, $4$ local chains, $2$ uniform chains and $5$ CGQeMCMC have found the minima. Considering the classical methods had an extra order of magnitude (due to computational limitations) to search the state space, it is clear that quantum proposal is considerably more efficient. It should be noted that to evaluate this system by brute force requires $\approx 3\times 10^{7}$ samples, meaning that the uniform proposal does not provide a practically useful proposal method, merely an upper bound on optimal thermalisation.
Figure 5: Cumulative magnetisation (top) of $10$ chains for uniform (orange), local (dark green) and CGQeMCMC (light green) with $q=\sqrt{n}=5=n_{g}$ ($5$ groups of $5$ spins) for $T=1$. The accuracy of average magnetism (bottom) compared to the actual value calculated by brute force erroneously suggests that the classical local has converged.
Analysis of the average cumulative magnetisation in Figure 5 incorrectly suggest that the classical local proposal converges particularly well, however further analysis (including Figure 4) suggests that local minima of similar magnetism to the global minima are contributing to no observable bias in the magnetisation. The CGQeMCMC does however observe a bias, due to some chains being stuck in local minima with different magnetisation. In the $25$ spin example, the $4^{th}$ excited state has a large Hamming distance from any other low energy state, being $12$ spin flips from any of the other lowest $10$ states. This is a very large Hamming distance, so traversing the energy landscape from the local minima is very difficult. Although any local proposal is likely to over-sample this region, the magnetisation of this state is the same as the ground state, meaning that this is not picked up in average magnetisation. This leads one to conclude that cumulative average magnetisation is not sufficient to fully characterise the thermalisation of a Markov chain.
@@
As only one model was considered for each system size and the average magnetisation provides biased results in this instance, no explicit results are reported from this section. It is clear however that even when only employing $5$ simulated qubits, a quantum enhanced search of the state space is more effective than both classical methods.
6 Conclusions
@@
In this paper starting with the QeMCMC approach of [1], we identified two important elements that lead to its desirable performance, namely the possibility to suggest proposal steps with big Hamming distance from the previous configuration (due to generating superposition of all configurations), while returning a new configuration that is close in energy with the previous one (due to the use of the problem Hamiltonian and the suitable annealing-style brief evolution). At the same time we identified, as a limitation of QeMCMC, the necessity for using QPUs of size as large as the Ising system considered.
@@
We then tested our proposal(s) using two figures of merit: The absolute spectral gap and magnetisation estimation. Spectral gap analysis of simulated quantum proposals in Section 5.1 was confined to systems comprising 10 spins or fewer. Nevertheless, the observed scalability of CGQeMCMC proposals relative to system size shows promising trends. Using only $\sqrt{n}$ qubits, Quantum-enhanced Markov Chain thermalisation is predicted to outperform classical techniques for larger systems, giving a quantum enhancement factor $k_{QEF}$ (see Section 5.1 and Equation 16 specifically) between $1.84$ and $2.6$.
The Markov Chain Monte Carlo algorithm was then executed in order to estimate the magnetisation for larger systems up to 36 spins, with coarse grained quantum proposals again using only $\sqrt{n}$ simulated qubits. Both classical proposals analysed here (local and uniform) struggle to reach thermalisation in these larger systems, falling into local minima or inefficiently exploring the large state-space. The coarse grained quantum proposals improve convergence overall for these larger systems, however they they also occasionally suffer from local minima.
There are many directions for future investigations related to the CGQeMCMC. The choice of $q=\sqrt{n}$ qubits in this paper was successful, however the optimal relation between $q$ and fastest thermalisation or indeed auto-correlation remains open. Likewise, the effect of noise has not been investigated. The QeMCMC method is by construction noise-resilient, and our CG approach uses smaller QPUs minimising further these effects, but still concrete analysis for the level and types of noise that can be accommodated maintaining the quantum advantage is another open question. Here it is also worth pointing out that due to the one-shot nature of the algorithm, standard error mitigation methods would fail and it would require a bespoke scheme that is outwith current literature [29, 30].
An important future step is to apply coarse graining to problems other than the Ising model. Application of variationally prepared Hamiltonian evolution is also an interesting avenue of research, as is coarse graining quantum algorithms that have already been derived from the QeMCMC [6, 7]. Other interesting next steps are to apply the CGQeMCMC to quantumly enhance classical methods such as simulated annealing, parallel tempering and population annealing.

### v2

@@ IV.1 Naive CG - local groups
Ignoring the surrounding spins is an approximation that understandably has limitations and specifically raises two main issues. Firstly, the Hamming distance of proposal steps is fundamentally restricted to the size of the group. While this is better than classical local move approaches, it still puts a bound on the jumps that the proposal steps can make, and is more likely to get stuck in local minima (as with the classical local proposal) and partially cancels one of the key properties that lead to the success of the QeMCMC - the large Hamming distance update proposals.
The second issue is that the new problem Hamiltonian $H_{prob}^{\prime}$ is missing contributions from the ignored spins making the evolution fundamentally incomplete. This partly cancels the other advantage of QeMCMC highlighted earlier, that it includes the problem’s Hamiltonian. If one were to consider a 1D Ising model where only neighboring spins interact, this choice of CG is likely to perform adequately, as the boundary interaction terms of the group is constant (2) with cluster size, meaning the impact of missing contributions from ignored spins scales favorably with group size. The same is not true for general Ising models with higher connectivity.
@@
IV.2 Improved local groups
@@
Explicitly, the interaction matrix is the same as in the local grouping, $\tilde{J}=J_{i\in g_{l},j\in g_{l}}$, but the non-selected spins are treated as environment, adding them to the field of the $j^{th}$ selected spin as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{h}_{j}=h_{j}+\sum_{i\notin g_{l}}^{n}J_{ji}s_{i}}$ (14)
Where each $s_{i}$ (value of spins not in the group) is now a constant parameter. The 2nd (field) term of the cost function now becomes:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{j\in g_{l}}\tilde{h}_{j}s_{j}=\sum_{j\in g_{l}}s_{j}(h_{j}+\sum_{i\notin g_{l}}J_{ji}s_{i})}$ (15)
In this way, the accuracy of the Hamiltonian is increased from the local group method without the addition of any quantum gates. The effective Hamiltonian $H_{prob}^{\prime}$ now fully represents the relative energy of any two states of the group. As will become clear in Section V, improved local grouping refines the naive local groups, however there is still tail-off for small $q/n$.
@@ IV.3 Multiple groups
// Initialization
2 $\mathbf{s}\leftarrow$ initial spin configuration;
@@ 8      $index\leftarrow\text{spins in }i^{th}\text{ group}$;
9      $s_{g}\leftarrow\text{subset of }\mathbf{s}\text{ using }index$;
@@ 10      $s_{g}^{\prime}\leftarrow$ QeMCMC ($s_{g}$);
      // update $\mathbf{s}^{\prime}$ with results from $i^{th}$ group’s QeMCMC
11      $\mathbf{s}^{\prime}[index]\leftarrow s_{g}^{\prime}$ ;
@@
V Numerical Results
In this section we analyse the results of our numerical experiments of Ising model CGeMCMC, while further details of the implementation can be found in Appendix A.
V.1 Spectral gap
To quantify the CGQeMCMC, the effect of reducing $q$ on the convergence of the resulting Markov chain must be understood, with the particular aim of understanding the different coarse grained proposals at $q=\sqrt{n}$. This is done by evaluating the spectral gap of the CG methods for randomly initialized fully-connected Ising models in the range $4-10$ spins. Figure 1 shows the decrease of spectral gap with $q/n$ for each of the CG methods and $n=\{6,8,10\}$. Clearly the decay is worse for techniques that do not multiple sample, suggesting that the locality of low $q$ restricts the proposal. This effect is dependent on system size, however it is difficult to quantify using only the small systems at our disposal. There is also a temperature dependence, as the performance of small $q$ is inextricably linked with the performance of the classical local proposal. For analysis of temperature dependence for a 9 spin instance see Figure 2.
@@ V.1 Spectral gap
In analogy with Layden et al. (2023) Figure 2a, the different update proposal strategies are compared for a representative set of $100$ randomly initialized $9$ spin fully-connected Ising model instances in Figure 2. The QeMCMC (blue) is far superior to any classical strategy (orange and green) in the low temperature regime. The CGQeMCMC with $3$ groups of $3$ qubits (red) is also superior to the classical strategy, however the speedup is clearly dampened. If one evaluates a single $3$ spin group on the lattice each quantum proposal (transparent blue) the quantum speedup is lost for most temperature ranges for this system size.
Figure 2: Spectral gap varying with temperature for $100$ randomly initialized $9$ spin fully-connected Ising model. The classical uniform (orange) and local (green) strategies are clearly improved at low temperatures by the QeMCMC strategy (blue) Layden et al. (2023). The red data represents CGQeMCMC with $q=\sqrt{n}=3=n_{g}$ ($3$ groups of $3$ spins). The light blue data is $q=\sqrt{n}$ but $n_{g}=1$ ($1$ group of $3$ spins).
There is evident promise of quantum speedup using merely $\sqrt{n}$ qubits, however speedup at a fixed system size is less important. The scaling behaviour of the proposal step is necessarily analysed at $T=1$ in Figure 3. Equation 16 is the exponential fitting function used in the figures, where a small $k$ represents good scaling of a proposal strategy. The data points for larger systems are less reliable due to a much restricted sample size $(50)$ for the average spectral gap. This leads to added uncertainty in the fits, and increased reliance on the convergence analysis in Section V.2.
@@
Figure 3: Spectral gap varying with problem size for $500$ randomly initialized fully-connected Ising models of each integer number of spins. Only $50$ systems of $9$ and $10$ spins for CG techniques were evaluated due to computational limitations. Each data set has least squares exponential fits with fitting function Equation 16. As the system size increases, the spectral gap of classical uniform (orange) and local (green) strategies clearly decrease faster than those of the QeMCMC (blue) and CGQeMCMC (red, light blue) proposals. For a given proposal, the value of $k$ determined by best fit is as follows: Uniform: $0.97(1)$, Local $0.92(4)$, QeMCMC: $0.335(8)$, local group: $0.77(2)$, Improved local group: $0.64(2)$, multiple sample: $0.50(3)$. $k$ values for $T=0.1$ and $T=10$ are listed in appendix 1 alongside the appropriate figures.
@@
For the multiple sample proposal strategy, we see a quantum enhancement factor of $k_{QEF}=1.84$ against the best classical approach - the local proposal. In this data, the QeMCMC gains an enhancement of $2.7$, a slightly reduced value than reported in Layden et al. (2023) for the same temperature. Figure 3 depends largely on the temperature while Figure 2 depends on the problem size. Appendix B provides multiple figures that explores these relations. The corresponding quantum speedups in $k$ for $T=0.1$ and $T=10$ are $2.17$ and $2.6$ respectively. Of course, the problem instance type (1D, 2D, fully connected, sparse etc.) influences the outcome of these experiments, however these topics are beyond the scope of this work.
@@ V.2 Convergence analysis
Markov chains using local, uniform and CGQeMCMC are used to evaluate the average energy and magnetization of the stationary distribution of exemplar fully-connected Ising models. Systems of size $16$, $25$ and $36$ (see Appendix B for $15$ and $25$ spin results) are the focus so that $\sqrt{n}$ qubits can be easily compared with classical methods. The stationary distribution of the $36$ spin instance at $T=0.1$ is the primary focus of this analysis as it is the largest system studied and has low temperature, which as seen in Figure 2 is where we expect CGQeMCMC to perform well. The problem is approaching that of an optimization problem (the limit reached at zero temperature), as most of the probability distribution lies in the ground state of the system, making it particularly difficult. See Appendix B for temperature and system size analysis.
@@
Figure 4: Average energy of $20$ chains at $100$ step intervals for uniform (orange), local(green), Improved local group without multiple sampling (light blue) and multiple sampling (red) with $q=\sqrt{n}=6=n_{g}$ ($6$ groups of $6$ spins) for $T=0.1$. Multiple sampling with $q>\sqrt{n}$ is also simulated, with 12 qubits (pink) and 18 qubits (purple). Right figure is close-up of the low energy region. The more local a proposal, the larger the offset from the true average energy (black line). This is greatly reduced by CGQeMCMC with multi-sampling and larger groups - methods that decrease the locality of proposals. The uniform proposal suffers from extremely slow convergence but does not get stuck in local minima.
@@
Figure 4 shows that all proposed methods are unable to converge within simulable timescales. The CGQeMCMC appears most effective, even though the problem of getting stuck in local minima persists since the average energy over $20$ Markov chains is biased by a few that have not yet “found” the minimum state. One can crudely evaluate convergence in this relatively low temperature regime by calculating how many chains have discovered the minimum state. In this example, where we consider $20$ chains for each proposal, $5$ local chains, 00 uniform chains and $11$ CGQeMCMC chains (6 qubit multiple sampling) have found the minima. The classical methods had an extra order of magnitude steps (due to computational limitations) to search the state space, and thus it is evident that quantum proposals are significantly more efficient. When one considers a Markov chain as part of a larger algorithm, with multiple parallel chains, this improvement in the quality of convergence becomes more substantial, even in the cases that it still fails to converge perfectly.
It should also be noted that increasing the number of qubits used (smaller coarse graining) has a positive effect on convergence. Simulations of 12 qubits (pink) and 18 qubits (purple) show increasing promise, despite being simulated for far fewer steps. This highlights the flexible nature of this coarse graining algorithm, that can take as input the size of the hardware that one has and accordingly achieve the best possible performance.
Figure 5: Average cumulative magnetization of $20$ chains of each proposal method where the colors are the same as in Figure 4. If the exact magnetization were not known, one may erroneously believe that some methods may have converged, when in reality none have.
The magnetization plots confirm lack of convergence of all proposed methods when compared with the brute force exact magnetization. It is important to note, that one cannot infer from the magnetization plots the convergence. Specifically, while a failure to converge in the magnetization plot also implies failure of convergence in general, the converse does not hold, since it is possible to have instances that appear to get the correct magnetization but fail to converge to the correct probability distribution. This becomes more important for “real life” examples where brute-force numerics are not possible.
@@
It is clear that even when employing only $6$ simulated qubits, a quantum enhanced exploration of the state space is more effective than both classical methods in the low temperature regime. One should note that when the temperature is raised to $T=1$, this quantum enhancement is diminished. See Appendix B for the complete set of plots for $T=\{0.1,1\}$ and $n=\{16,25,36\}$.
Although the results reported in this section point towards to a scaling advantage, this cannot be explicitly proven by analysis of only a select few Ising model instances. A more thorough analysis with multiple different models would be required.
VI Conclusions
@@
In this paper starting with the QeMCMC approach of Layden et al. (2023), we identified two important elements that lead to its desirable performance, namely the ability to suggest proposal steps with big Hamming distance from the previous configuration (due to generating superposition of all configurations), while returning a new configuration that is close in energy with the previous one (due to the use of the problem Hamiltonian and the suitable annealing-style brief evolution). At the same time, we identified a limitation of QeMCMC regarding the requirement of a QPU with as many qubits as the size of the Ising system considered.
@@
We then tested our proposal(s) using two figures of merit: The absolute spectral gap and magnetization estimation. Spectral gap analysis of simulated quantum proposals in Section V.1 was confined to systems comprising 10 spins or fewer. Nevertheless, the observed scalability of CGQeMCMC proposals relative to system size shows promising trends. Using only $\sqrt{n}$ qubits, Quantum-enhanced Markov Chain thermalization is predicted to outperform classical techniques for larger systems, giving a quantum enhancement factor $k_{QEF}$ (see Section V.1 and Equation 17 specifically) between $1.84$ and $2.6$.
The Markov Chain Monte Carlo algorithm was then executed in order to estimate the magnetization for larger systems up to 36 spins, with coarse grained quantum proposals again using only $\sqrt{n}$ simulated qubits. Both classical proposals analysed here (local and uniform) struggle to reach thermalization in these larger systems, falling into local minima or inefficiently exploring the large state-space. The coarse grained quantum proposals improve convergence overall for these larger systems, however they they also occasionally suffer from local minima.
There are many directions for future investigations related to the CGQeMCMC. The choice of $q=\sqrt{n}$ qubits in this paper was successful, however the optimal relation between $q$ and fastest thermalization or indeed auto-correlation is both problem and hardware dependent. Future work should use the CGQeMCMC to exploit the connectivity of hardware. For example, if one were to consider an Ising model with 3D connectivity, they could partition the problem into $n$ planar lattices that are suitable for the 2D connectivity common to many QPU. In this way, the CGQeMCMC can be a powerful tool in relating highly connected problems to QPUs with limited connectivity.
The effect of noise has not been investigated. The QeMCMC method is by construction noise-resilient, and our CG approach uses smaller QPUs minimizing further these effects. Concrete analysis for the level and types of noise under which we maintain the quantum scaling advantage is an important open question. Here it is also worth pointing out that due to the one-shot nature of the algorithm, standard error mitigation methods are not applicable meaning it would require a bespoke scheme that is outwith current literature Li and Benjamin (2017); Temme et al. (2017).
Another future direction is quantification of the sampling error induced by coarse graining. A Markov chain produced by the CGQeMCMC provably samples the Boltzmann distribution as it satisfies detailed balance and is aperiodic and irreducible. The number of steps required to produce enough samples to estimate some observable within a given error is dependent on the auto-correlation time of the chain. Speeding up thermalization does not necessarily result in quicker auto-correlation, and thus this must be studied in future work to properly understand the cost of coarse graining compared to the QeMCMC.
An important future step is to apply coarse graining to problems other than the Ising model. Application of variationally prepared Hamiltonian evolution is also an interesting avenue of research, as is coarse graining quantum algorithms that have already been derived from the QeMCMC Lockwood et al. (2024); Nakano et al. (2023). Other interesting next steps are to apply the CGQeMCMC to quantumly enhance classical methods such as simulated annealing, parallel tempering and population annealing.